what should be subtracted from twice the rational number -3/11 to get -5/22?
Answer by fractalier(6550)

We can write this as
2(-3/11) - x = -5/22
and figure it as
-6/11 - x = -5/22
Let us multiply by -22 and get
12 + 22x = 5
22x = -7
x = -7/22